Find the differential equation to the following;

dy/dx=9.8-0.196y


1
Expert's answer
2021-07-15T13:13:22-0400

Let us solve the differential equation "\frac{dy}{dx}=9.8-0.196y." It follows that "\frac{dy}{9.8-0.196y}=dx," and hence "\int\frac{dy}{9.8-0.196y}=\int dx." Then "-\frac{1}{0.196}\int\frac{d(9.8-0.196y)}{9.8-0.196y}=\int dx," and thus "-\frac{1}{0.196}\ln|9.8-0.196y|=x+C". It follows that "\ln|9.8-0.196y|=-0.196x+C_1" or "9.8-0.196y=C_2e^{-0.196x}." We conclude that the general solution is of the form "y=50-C_3e^{-0.196x}."